OpenStudy (anonymous):

What two numbers whose product is -12 and sum of 9

hero (hero):

Hint: ab = -12 a + b = 9

OpenStudy (anonymous):

Yes I no that but I can't seem to find the two number

hero (hero):

2nd Hint: a = 9 - b

OpenStudy (anonymous):

How do you find b

hero (hero):

b = -12/a

OpenStudy (anonymous):

-21 and 9?

hero (hero):

Where'd you get those numbers from?

OpenStudy (anonymous):

I did something wrong

OpenStudy (anonymous):

I don't no how to do this

hero (hero):

Well, that's because you don't know the secret trick.

OpenStudy (anonymous):

And that is

hero (hero):

ab = -12 a + b = 9 Substitute a into the first equation to get \((9-b)b = -12 \\ -b^2 + 9b = -12 \\ -(b^2 - 9b) = -12 \\ b^2 - 9b = 12 \\ b^2 - 9b - 12 = 0 \)

hero (hero):

It's a quadratic equation

OpenStudy (anonymous):

Sou solve from there ?

hero (hero):

Yes. You might want to try completing the square for this.

hero (hero):

\(b^2 - 9b = 12 \)

OpenStudy (anonymous):

How do you solve to get the two numbers from here?

hero (hero):

Add 81/4 to both sides \(b^2 - 9b + \frac{81}{4} = 12 + \frac{81}{4}\)

OpenStudy (anonymous):

How did you get 81/4

OpenStudy (anonymous):

you dontt even need to do all that. just make a factor tree for -12 and see what two pairs add to 9

hero (hero):

I got 81/4 by dividing the coefficient of the middle term by 2, then squaring it. Simplify the left side into a binomial: \((b - \frac{9}{2})^2 = 12 + \frac{81}{4}\)

hero (hero):

Taking the square root of both sides to get \(b - \frac{9}{2} = \sqrt{12 + \frac{81}{4}}\)

hero (hero):

Then adding 9/2 to both sides to get: \(b = \sqrt{12 + \frac{81}{4}} + \frac{9}{2}\)

hero (hero):

Actually it's supposed to be + or -

hero (hero):

\(b = \pm \sqrt{12 + \frac{81}{4}} + \frac{9}{2}\)

hero (hero):

Now all we have to do is simplify that

hero (hero):

\(b = \pm \sqrt{\frac{48}{4}+ \frac{81}{4}} + \frac{9}{2}\)

hero (hero):

\(b = \pm \sqrt{\frac{48 + 81}{4}} + \frac{9}{2}\)

hero (hero):

\(b = \pm \sqrt{\frac{129}{4}} + \frac{9}{2}\)

hero (hero):

\(b = \pm \frac{\sqrt{129}}{\sqrt{4}} + \frac{9}{2}\)

hero (hero):

\(b = \pm \frac{\sqrt{129}}{2} + \frac{9}{2}\)

hero (hero):

\(b = \frac{\sqrt{129} + 9}{2}\)

hero (hero):

Good luck finding a

hero (hero):

Better yet use a = 9 - b to find a

hero (hero):

\(a = 9 - \frac{\sqrt{129} + 9}{2}\)

hero (hero):

\(a = \frac{18}{2} - \frac{\sqrt{129} + 9}{2}\)

mathslover (mathslover):

@Hero ab = -12 a + b = 9 \[\large{\color{blue}{(a+b)^2=81}}\] \[\large{\color{red}{a^2+b^2+2ab=81}}\] \[\large{\color{magenta}{a^2+b^2+2(-12)=81}}\] \[\large{\color{brown}{a^2+b^2=81+24}}\] \[\large{\color{green}{a^2+b^2=105}}\] \[\large{\color{orange}{a^2-b^2=\sqrt{(a^2+b^2)^2-4a^2b^2}}}\] \[\large{\color{blue}{a^2-b^2=\sqrt{(105)^2-(144*4)}}}\] \[\large{\color{brown}{a^2-b^2=102->\textbf{Approx.}}}\] \[\large{\color{blue}{a^2+b^2=105}}\] \[\large{\color{red}{a^2-b^2=102}}\] \[\large{\color{orange}{a^2=105-b^2}}\] \[\large{\color{brown}{\textbf{Put this value of a^2 in the second equation of a^2-b^2}}}\] \[\large{\color{red}{105-b^2-b^2=102}}\] \[\large{\color{orange}{-2b^2=-207}}\] \[\large{\color{magenta}{b^2=\frac{-207}{-2}}}\] \[\large{\color{blue}{b=\sqrt{\frac{207}{2}}}}\] \[\large{\color{green}{b=10.17}}\] \[\large{\color{red}{a=9-10.17=-1.17}}\]

mathslover (mathslover):

I don't know whether i am right or wrong..

hero (hero):

Yeah, but I still have exact vaules

mathslover (mathslover):

Yes you have them i am just off with .11 for product

hero (hero):

Nice colors

mathslover (mathslover):

thanks... well yes your's one gives b = 10.17 but a is giving something wrong

hero (hero):

Well, both a and b can be either positive or negative

mathslover (mathslover):

but a is something bigger than 12 at least... sorry got to go for school now... if i would not then there will be a stick ready for my body .. :)

mathslover (mathslover):

given ab = -12 but product of -19.17 and 10.17 is not near to 12

mathslover (mathslover):

*-12

hero (hero):

I know what I did wrong now. \(a = \frac{18 - \sqrt{129} - 9}{2}\)

OpenStudy (anonymous):

wow

hero (hero):

lol

OpenStudy (anonymous):

what is wrong with the quadratic formula?

OpenStudy (anonymous):

\(x^2+9x-12=0\) etc

hero (hero):

Nothing.

hero (hero):

Well...nobody likes using it

OpenStudy (anonymous):

You cant use factoring for this?

hero (hero):

Therefore, a = \(\large\frac{9 - \sqrt{129}}{2}\) b = \(\large\frac{( 9+ \sqrt{129}}{2}\)

hero (hero):

You can use factoring, quadratic formula, or complete the square for any quadratic.